# Vegetation Patterning Can Both Impede and Trigger Critical Transitions from Savanna to Grassland

### v1

@@ 1 Introduction
The coexistence of trees and C4 grasses defines savannas, which support a significant proportion of the global biodiversity and human populations (Sankaran et al., 2005; Parr et al., 2014). The long-standing question of how trees and grasses persist together, despite competing for the same limited resources like water, is essential for assessing the existence, stability and resilience to environmental change of savanna ecosystems. A recent review (Holdo and Nippert, 2023) clarifies the current understanding of the coexistence problem, distinguishing between mesic savannas (mean annual precipitation $>650$ mm) and dry savannas ($<650$ mm) (Sankaran et al., 2005; D’Onofrio et al., 2018; Holdo et al., 2018). In mesic savannas, tree-grass balance is maintained primarily by disturbances such as herbivory and fire, while in dry savannas, water stress plays a dominant role. Despite these advances, recent models overlook key dynamics specific to dry savannas, particularly the nature of tree-grass interactions including their spatially extended structure.
@@
Vegetation patterns are often viewed as early warning signals of ecological tipping points that may lead to ecosystem collapse (Rietkerk et al., 2004). Simultaneously, pattern formation may also enhance ecosystem resilience by delaying the onset of such tipping points (Rietkerk et al., 2021). For instance, spatially explicit models reveal that patterns can remain stable even under parameter changes that would induce collapse in corresponding non-spatial models (Lefever and Lejeune, 1997; von Hardenberg et al., 2001; Rietkerk et al., 2002). This phenomenon, which we refer to as “Turing-evades-tipping” (Rietkerk et al., 2021), emphasizes the the significance of spatial interactions in enhancing ecosystem resilience and underscores the limitations of non-spatial approaches (Banerjee et al., 2024). Various spatial models have successfully captured vegetation patterning and pattern dynamics across diverse ecosystems (Klausmeier, 1999; Hille Ris Lambers et al., 2001; Rietkerk et al., 2002; Gilad et al., 2007b; Zelnik et al., 2015; Getzin et al., 2016; Bastiaansen et al., 2018; Bennett et al., 2023a, b), including regular tree-grass patterning in a savanna environment (Tega et al., 2022) and shifts from competition to facilitation consistent with the stress-gradient hypothesis (Gilad et al., 2007a). However, these models generally lack life-stage specific dynamics, a critical component in savanna ecosystems.
@@ 2 Methods
$a_{g}$ Growth rate of grass $1.5-5.1$ yr-1
$K_{g}$ Theoretical grass biomass cap $0.2-2.0$ kg m-2
$m_{g}$ Mortality rate of grass $1.5-4.6$ yr-1
$p_{f}$ Rate of facilitation $0-2$ yr-1
$d_{g}$ Diffusion rate of grass $50-100$ m2 yr-1
$a_{s}$ Growth rate of trees $0.001-0.1$ yr-1
$C_{s}$ Threshold tree density $1000-K_{s}/$ 100 kg m-2
$K_{s}$ Theoretical tree biomass cap $3-10$ kg m-2
$m_{s}$ Mortality rate of trees $0.046-0.46$ yr-1
$p_{c}$ Rate of competition $5.99-34.5$ yr-1
$d_{s}$ Diffusion rate of trees $5-10$ m2 yr-1
@@
Grass growth is modelled using a logistic function, reflecting how grasses expand toward a maximum standing biomass, typical of the continuous grass canopy in dry savannas despite resource constraints like water (Holdo and Nippert, 2023). Assuming $a_{g}>m_{g}$ excludes bare soil stability, aligning with a focus on savanna ecosystems where grasses are relatively abundant. The model can be adapted in a straightforward manner to include bare soil dynamics by replacing the logistic growth function with an Allee-effect growth function, mimicking vegetation-water feedbacks observed in existing dryland models (Klausmeier, 1999; Rietkerk et al., 2002; Bennett et al., 2023a).
@@
In their early-life stages, savanna trees establish roots in the upper soil layers where grasses limit water access for juvenile trees, inhibiting their growth and establishment (Holdo and Nippert, 2023). This competitive inhibition is modelled by the term $-p_{c}gs\beta(s)$, which represents an additional mortality factor at low tree biomass densities, reflecting the intense competition young trees face from grasses. As trees mature and develop deeper roots, they access water unavailable to grasses, reducing this competitive pressure. The function $\beta(s)$ captures this shift, inducing that grass competition remains strong at low tree biomass but decreases as tree biomass increases (Figure 2).
@@
We model the spatial effects of seed dispersal using diffusion, a standard approach in spatially extended vegetation models (Klausmeier, 1999; Hille Ris Lambers et al., 2001; Gilad et al., 2007b; Bennett et al., 2023b). In dry savannas, grasses propagate rapidly, forming a nearly continuous grass layer (Holdo and Nippert, 2023). In contrast, savanna trees do not form closed canopies in these arid environments (Sankaran et al., 2005). Although tree seed dispersal rates may be comparable to those of grasses, tree seeds face greater challenges in germination, and newly emerged seedlings require favorable conditions to survive, resulting in a lower overall rate of biomass redistribution. To capture these dynamics, we assign a higher diffusion coefficient to grasses ($d_{g}$) than to trees ($d_{s}$), as the diffusion coefficient reflects not only the rate of spatial seed spread but also the constraints on the redistribution of biomass imposed by germination and establishment. When spatial effects are excluded, i.e., when $d_{g}=d_{s}=0$, the model simplifies to a non-spatial system. This non-spatial version is useful for studying the coexistence problem and provides a baseline for comparison. We refer to the spatially extended system as the “spatial model” and the version without diffusion as the “non-spatial model”.
@@ 4 Discussion
Our model captures the key interactions driving tree-grass dynamics in dry savannas, as outlined by (Holdo and Nippert, 2023). Most of these mechanisms are water-mediated, and future work could include water as an explicit variable for a more realistic and mechanistic modelling effort, which could possibly be better parameterized from observations. This would offer a clearer understanding of how water availability influences ecosystem stability and collapse, addressing potential limitations of our simplified model. Moreover, incorporating other ecological factors, such as fire regimes and herbivory (more critical in mesic savannas (Sankaran et al., 2005; D’Onofrio et al., 2018)), adding complex non-local seed dispersal mechanisms (Patterson et al., 2024) or considering phenotypic transitions (Bennett et al., 2023b), would further enhance our understanding of how spatial patterning impacts long-term savanna dynamics.
@@ Appendix A Parametrization and calibration
  • $a_{g}$ (Growth rate of grass): This parameter governs the exponential growth rate of grasses in absence of trees and far away from the maximum standing biomass. Data suggests that the relative growth rate of grasses can reach values up to $0.3$ yr-1 (Simpson, 2023). To account for variability, we set an upper bound of $0.5$, resulting in $1.51<e^{a_{g}-m_{g}}<$ 1.5. Thus, we find $m_{g}<a_{g}<m_{g}+\ln{1.5}$ yr-1.
  • $K_{g}$ (Theoretical grass biomass cap): According to (Moustakas et al., 2013), grass biomass densities further away from the tree canopy in drier sites typically range between $0.2$ and $0.5$ kg m-2. Based on this observation, we impose the condition $0.50.2<g_{cc}<$ 0.5 kg m-2, where $g_{cc}=K_{g}\left(1-\frac{m_{g}}{a_{g}}\right)$ represents the grass carrying capacity in absence of trees. Considering the parameter ranges for $a_{g}$ and $m_{g}$, we found the range $20.2<K_{g}<$ 2 kg m-2 to adequately capture this constraint.
  • $K_{s}$ (Theoretical tree biomass cap): Based on biomass estimates of savanna trees across precipitation gradients in dry savannas, tree biomass densities have been estimated as $2.8$ kg m-2 in a tree savanna and as $5$ kg m-2 in an open woodland (Food and Agriculture Organization of the United Nations, 1997). Therefore, we impose the condition $52.8<\tilde{s}<$ 5 kg m-2, where $\tilde{s}$ represents the carrying capacity for savanna trees in absence of grasses, as introduced in Appendix B. Since $K_{s}$ represents an upper biomass limit, we extend the upper bound and define the range as $102.8<K_{s}<$ 10 kg m-2.
@@
  • $p_{c}$ (Rate of competition): When no mature trees are present, $s$ is small, so $\beta(s)\approx 1$. At the grass carrying capacity $g_{cc}$, competition reduces to $-p_{c}g_{cc}s$. Early-life stage trees have very low survival rates due to intense competition with grasses (Sankaran et al., 2004; Riginos, 2009), so we assume $95$ to $99.9$ perish within a year. This gives the bounds $\ln(20)/g_{cc}<p_{c}<\ln(1000)/g_{cc}$ which simplifies to $2\ln(20)<p_{c}<5\ln(1000)$ based on the extremes of $g_{cc}$.
@@
  • $s_{c}$ (Characteristic parameter of $\beta(s)$): At the tree carrying capacity ($s=\tilde{s}$), $\beta(s)$ should be relatively close to 00. Quantifying “close”, we set this to at most $20$ percent of $1$, yielding the range given by $s_{c}=\tilde{s}\cdot\left(\frac{C}{1-C}\right)^{\frac{1}{n}}$ for $C\in(0,0.2]$.
@@
  • $a_{s}$ (Growth rate of trees): This parameter controls savanna tree growth. Over the gradient of this parameter we expect to see the transition between savanna and grassland. Calibration yields the range $0.10.001<a_{s}<$ 0.1 yr-1.
  • $p_{f}$ (Rate of facilitation): This parameter strongly influences the increase in grass biomass under the tree canopy due to shading effects. Based on (Moustakas et al., 2013), this facilitation should not exceed $100$ percent. Model calibration suggest the range $20<p_{f}<$ 2 yr-1.
  • $d_{g}$ (Diffusion rates of grasses): Based on model calibration, we set $10010<d_{g}<$ 100 m2 yr-1, ensuring realistic pattern wavelengths for a $400$ m x $400$ m spatial grid, in line with observed scales of patterning (Douma et al., 2007).
@@ Appendix B Linear stability analysis
  • Tree-only states: $(0,\tilde{s})$, where $\tilde{s}$ is a solution to the quadratic equation:

### v2

@@ 1 Introduction
The coexistence of trees and C4 grasses defines savannas, which support a significant proportion of the global biodiversity and human populations (Sankaran et al., 2005; Parr et al., 2014). The long-standing question of how trees and grasses persist together, despite competing for the same limited resources like water, is essential for assessing the existence, stability and resilience to environmental change of savanna ecosystems. A recent review (Holdo and Nippert, 2023) clarifies the current understanding of the coexistence problem, distinguishing between mesic savannas (mean annual precipitation >650>650 mm) and dry savannas (<650<650 mm) (Sankaran et al., 2005; D’Onofrio et al., 2018; Holdo et al., 2018). In mesic savannas, tree-grass balance is maintained primarily by disturbances such as herbivory and fire, while in dry savannas, water stress plays a dominant role. Despite these advances, recent models overlook key dynamics specific to dry savannas, particularly the nature of tree-grass interactions including their spatially extended structure.
@@
Vegetation patterns are often viewed as early warning signals of ecological tipping points that may lead to ecosystem collapse (Rietkerk et al., 2004). Simultaneously, pattern formation may also enhance ecosystem resilience by delaying the onset of such tipping points (Rietkerk et al., 2021). For instance, spatially explicit models reveal that patterns can remain stable even under parameter changes that would induce collapse in corresponding non-spatial models (Lefever and Lejeune, 1997; von Hardenberg et al., 2001; Rietkerk et al., 2002). This phenomenon, which we refer to as “Turing-evades-tipping” (Rietkerk et al., 2021), emphasizes the the significance of spatial interactions in enhancing ecosystem resilience and underscores the limitations of non-spatial approaches (Banerjee et al., 2024). Various spatial models have successfully captured vegetation patterning and pattern dynamics across diverse ecosystems (Klausmeier, 1999; Hille Ris Lambers et al., 2001; Rietkerk et al., 2002; Gilad et al., 2007b; Zelnik et al., 2015; Getzin et al., 2016; Bastiaansen et al., 2018; Bennett et al., 2023b, a), including regular tree-grass patterning in a savanna environment (Tega et al., 2022) and shifts from competition to facilitation consistent with the stress-gradient hypothesis (Gilad et al., 2007a). However, these models generally lack life-stage specific dynamics, a critical component in savanna ecosystems.
@@ 2 Methods
$a_{g}$ Growth rate of grass 1.55.11.5-5.1 yr-1
$K_{g}$ Theoretical grass biomass cap 0.22.00.2-2.0 kg m-2
$m_{g}$ Mortality rate of grass 1.54.61.5-4.6 yr-1
$p_{f}$ Rate of facilitation 020-2 yr-1
$d_{g}$ Diffusion rate of grass 5010050-100 m2 yr-1
$a_{s}$ Growth rate of trees 0.0010.10.001-0.1 yr-1
$C_{s}$ Threshold tree density $0-K_{s}/100$ kg m-2
$K_{s}$ Theoretical tree biomass cap 3103-10 kg m-2
$m_{s}$ Mortality rate of trees 0.0460.460.046-0.46 yr-1
$p_{c}$ Rate of competition 5.9934.55.99-34.5 yr-1
$d_{s}$ Diffusion rate of trees 5105-10 m2 yr-1
@@
Grass growth is modelled using a logistic function, reflecting how grasses expand toward a maximum standing biomass, typical of the continuous grass canopy in dry savannas despite resource constraints like water (Holdo and Nippert, 2023). Assuming $a_{g}>m_{g}$ excludes bare soil stability, aligning with a focus on savanna ecosystems where grasses are relatively abundant. The model can be adapted in a straightforward manner to include bare soil dynamics by replacing the logistic growth function with an Allee-effect growth function, mimicking vegetation-water feedbacks observed in existing dryland models (Klausmeier, 1999; Rietkerk et al., 2002; Bennett et al., 2023b).
@@
In their early-life stages, savanna trees establish roots in the upper soil layers where grasses limit water access for juvenile trees, inhibiting their growth and establishment (Holdo and Nippert, 2023). This competitive inhibition is modelled by the term $p_{c}gs\beta(s)$, which represents an additional mortality factor at low tree biomass densities, reflecting the intense competition young trees face from grasses. As trees mature and develop deeper roots, they access water unavailable to grasses, reducing this competitive pressure. The function $\beta(s)$ captures this shift, inducing that grass competition remains strong at low tree biomass but decreases as tree biomass increases (Figure 2).
@@
We model the spatial effects of seed dispersal using diffusion, a standard approach in spatially extended vegetation models (Klausmeier, 1999; Hille Ris Lambers et al., 2001; Gilad et al., 2007b; Bennett et al., 2023a). In dry savannas, grasses propagate rapidly, forming a nearly continuous grass layer (Holdo and Nippert, 2023). In contrast, savanna trees do not form closed canopies in these arid environments (Sankaran et al., 2005). Although tree seed dispersal rates may be comparable to those of grasses, tree seeds face greater challenges in germination, and newly emerged seedlings require favorable conditions to survive, resulting in a lower overall rate of biomass redistribution. To capture these dynamics, we assign a higher diffusion coefficient to grasses ($d_{g}$) than to trees ($d_{s}$), as the diffusion coefficient reflects not only the rate of spatial seed spread but also the constraints on the redistribution of biomass imposed by germination and establishment. When spatial effects are excluded, i.e., when $d_{g}=d_{s}=0$, the model simplifies to a non-spatial system. This non-spatial version is useful for studying the coexistence problem and provides a baseline for comparison. We refer to the spatially extended system as the “spatial model” and the version without diffusion as the “non-spatial model”.
@@ 4 Discussion
Our model captures the key interactions driving tree-grass dynamics in dry savannas, as outlined by (Holdo and Nippert, 2023). Most of these mechanisms are water-mediated, and future work could include water as an explicit variable for a more realistic and mechanistic modelling effort, which could possibly be better parameterized from observations. This would offer a clearer understanding of how water availability influences ecosystem stability and collapse, addressing potential limitations of our simplified model. Moreover, incorporating other ecological factors, such as fire regimes and herbivory (more critical in mesic savannas (Sankaran et al., 2005; D’Onofrio et al., 2018)), adding complex non-local seed dispersal mechanisms (Patterson et al., 2024) or considering phenotypic transitions (Bennett et al., 2023a), would further enhance our understanding of how spatial patterning impacts long-term savanna dynamics.
@@ Appendix A Parametrization and calibration
  • $a_{g}$ (Growth rate of grass): This parameter governs the exponential growth rate of grasses in absence of trees and far away from the maximum standing biomass. Data suggests that the relative growth rate of grasses can reach values up to $0.3$ yr-1 (Simpson, 2023). To account for variability, we set an upper bound of $0.5$, resulting in $1<e^{a_{g}-m_{g}}<1.5$. Thus, we find $m_{g}<a_{g}<m_{g}+\ln{1.5}$ yr-1.
  • $K_{g}$ (Theoretical grass biomass cap): According to (Moustakas et al., 2013), grass biomass densities further away from the tree canopy in drier sites typically range between $0.2$ and $0.5$ kg m-2. Based on this observation, we impose the condition $0.2<g_{cc}<0.5$ kg m-2, where $g_{cc}=K_{g}\left(1-\frac{m_{g}}{a_{g}}\right)$ represents the grass carrying capacity in absence of trees. Considering the parameter ranges for $a_{g}$ and $m_{g}$, we found the range $0.2<K_{g}<2$ kg m-2 to adequately capture this constraint.
  • $K_{s}$ (Theoretical tree biomass cap): Based on biomass estimates of savanna trees across precipitation gradients in dry savannas, tree biomass densities have been estimated as $2.8$ kg m-2 in a tree savanna and as $5$ kg m-2 in an open woodland (Food and Agriculture Organization of the United Nations, 1997). Therefore, we impose the condition $2.8<\tilde{s}<5$ kg m-2, where $\tilde{s}$ represents the carrying capacity for savanna trees in absence of grasses, as introduced in Appendix B. Since $K_{s}$ represents an upper biomass limit, we extend the upper bound and define the range as $2.8<K_{s}<10$ kg m-2.
@@
  • $p_{c}$ (Rate of competition): When no mature trees are present, $s$ is small, so $\beta(s)\approx 1$. At the grass carrying capacity $g_{cc}$, competition reduces to $p_{c}g_{cc}s$. Early-life stage trees have very low survival rates due to intense competition with grasses (Sankaran et al., 2004; Riginos, 2009), so we assume $95$ to $99.9$ perish within a year. This gives the bounds $\ln(20)/g_{cc}<p_{c}<\ln(1000)/g_{cc}$ which simplifies to $2\ln(20)<p_{c}<5\ln(1000)$ based on the extremes of $g_{cc}$.
@@
  • $s_{c}$ (Characteristic parameter of $\beta(s)$): At the tree carrying capacity ($s=\tilde{s}$), $\beta(s)$ should be relatively close to 0. Quantifying “close”, we set this to at most $20$ percent of $1$, yielding the range given by $s_{c}=\tilde{s}\cdot\left(\frac{C}{1-C}\right)^{\frac{1}{n}}$ for $C\in(0,0.2]$.
@@
  • $a_{s}$ (Growth rate of trees): This parameter controls savanna tree growth. Over the gradient of this parameter we expect to see the transition between savanna and grassland. Calibration yields the range $0.001<a_{s}<0.1$ yr-1.
  • $p_{f}$ (Rate of facilitation): This parameter strongly influences the increase in grass biomass under the tree canopy due to shading effects. Based on (Moustakas et al., 2013), this facilitation should not exceed $100$ percent. Model calibration suggest the range $0<p_{f}<2$ yr-1.
  • $d_{g}$ (Diffusion rates of grasses): Based on model calibration, we set $10<d_{g}<100$ m2 yr-1, ensuring realistic pattern wavelengths for a $400$ m x $400$ m spatial grid, in line with observed scales of patterning (Douma et al., 2007).
@@ Appendix B Linear stability analysis
[…]